OpenStudy (amistre64):

Something neat I read, but not quite understand: Spose you have a sequence and need to find its generating function: \[a_1,a_2,a_3,a_4,a_5,...\] find the differences between \(a_n\) and \(a_{n+1\), such that they are \[b_1,b_2,b_3,b_4,b_5,...\] continue finding differences till you end up with a constant difference: \[k,k,k,k,k....\] the generating function can be determined by: \[a_1*n\choose 0 + b_1*n\choose 1 + c_1*n\choose 2 + ... + k* n\choose kth.row\] maybe .... Something neat I read, but not quite understand: Spose you have a sequence and need to find its generating function: \[a_1,a_2,a_3,a_4,a_5,...\] find the differences between \(a_n\) and \(a_{n+1\), such that they are \[b_1,b_2,b_3,b_4,b_5,...\] continue finding differences till you end up with a constant difference: \[k,k,k,k,k....\] the generating function can be determined by: \[a_1*n\choose 0 + b_1*n\choose 1 + c_1*n\choose 2 + ... + k* n\choose kth.row\] maybe .... @Mathematics

OpenStudy (amistre64):

ack!! ... latex failed me lol

OpenStudy (anonymous):

Yes I think that would work.

OpenStudy (amistre64):

Something neat I read, but not quite understand: Spose you have a sequence and need to find its generating function: \[a_1,a_2,a_3,a_4,a_5,...\] find the differences between \(a_n\) and \(a_{n+1}\), such that they are \[b_1,b_2,b_3,b_4,b_5,...\] continue finding differences till you end up with a constant difference: \[k,k,k,k,k....\] the generating function can be determined by: \[ a_1 {n\choose 0} + b_1 {n\choose 1}+ c_1{n\choose 2} + ... + k{ n\choose kth.row}\] maybe ....

OpenStudy (amistre64):

for example the sequence: 1, 4, 10, 23, 47, ... r0: 1 4 10 23 47 ... r1: 3 6 13 24 r2: 3 7 11 r3: 4 4 \[1{n\choose0}+3{n\choose1}+3{n\choose2}+4{n\choose3}\] \[1 + 3n + \frac{3}{2}n(n-1)+\frac{4}{6}n(n-1)(n-2)\]

OpenStudy (amistre64):

n=0,1,2,3,... of course

OpenStudy (amistre64):

its reminiscent of a taylor series to me .... almost

OpenStudy (anonymous):

so i have a question. suppose we look at the simplest case 1, 2, 3, 4, 5, ... what do we get?

OpenStudy (anonymous):

or can we use it to find summation as in \[\sum_{k=1}^nk\]

OpenStudy (amistre64):

1 2 3 4 5 1 1 1 1 \[1{n\choose0}+1{n\choose1}=1+n\] for the function that generates terms, n=0,1,2,3,...

OpenStudy (anonymous):

r0 1 3 6 10 15 21 2 3 4 5 6 1 1 1 1

OpenStudy (anonymous):

I think the way this works is as follows. Say I have a, a+d, a+2d, a+3d, ... (an arithmetic progression) The generating function is obviously a+(n-1)d. The difference of each term from the previous one is always d. Now suppose I have a, a+(a+d), a+(a+d)+(a+2d), a+(a+d)+(a+2d)+(a+3d), ..., where the difference of the terms form an arithmetic progression. This is simplified to a, 2a+d, 3a+3d, 4a+6d, ... The difference between a_n and a_{n-1} now is a+(n-1)d. Hence the generating function is a + summation of all these (n-1) a+(n-1)d's, which is a + [summation of (n-1) a's] + [summation of (n-1) (n-1)d's] The summation of (n-1) a's is a(n-1). The summation of (n-1) (n-1)d's is d(n-1)n/2. Hence general term is a + (n-1)a + d(n-1)n/2 = an + dn(n-1)/2. I think that if this argument is continued, the result you mention will follow inductively.

OpenStudy (anonymous):

so let me see if i understand. it should be, for the one i wrote \[1\dbinom{n}{0}+2\dbinom{n}{1}+\dbinom{n}{2}\]

OpenStudy (amistre64):

\[1{n\choose0}+2{n\choose1}+1{n\choose2}\] \[1+2n+\frac{n(n-1)}{2}=\frac{2+4n+n^2-n}{2}\] \[\frac{2+4n+n^2-n}{2}=\frac{1}{2}(n^2+3n+2)\]

OpenStudy (amistre64):

n=0, a0=1 n=1, a1=3

OpenStudy (amistre64):

good explanation xact

OpenStudy (amistre64):

\[\frac{1}{2}(5^2+3.5+2)=\frac{1}{2}(42)=21\]

OpenStudy (anonymous):

yes a good explanation. but i seem to have confused myself. why don't i just get \[\frac{n(n+1)}{2}\] instead of \[\frac{(n+1)(n+2)}{2}\]or am i just being dense?

OpenStudy (amistre64):

i think its the result of n=0,1,2,3,... as the domain

OpenStudy (anonymous):

ooooh starting at n = 0 i see!

OpenStudy (anonymous):

how about fibonacci??

OpenStudy (amistre64):

if the fib can get differences down to a constant ... might have to start at something other than the first term maybe

OpenStudy (anonymous):

ok i see this is "if you can do this" got it. you cannot for fibonacci because successive differences are fibonacci again, all the way down the line

OpenStudy (anonymous):

i have read maybe 20 times how to get generating functions for sequences and even gave a brief lesson on it, but for some reason it does not stick. to get the one for fib you solve a basic quadratic equation, but i forget the procedure

OpenStudy (amistre64):

i cant recall fibs either ;)